# FineReason: Evaluating and Improving LLMs’ Deliberate Reasoning through Reflective Puzzle Solving

### v1

@@ 4.4 Quality Analysis of State Checking
To examine the errors made in state checking, we conduct a human analysis of the mistakes from the best-performing model, o1, in Figure 3. The most common error is Misinterpretation of Premises, where o1 incorrectly uses available information to reach a faulty conclusion. For instance, in a grid puzzle, given the clue “The chocolate piece, Joey’s cake, and the \$125 cake are three different cakes,” the model still mistakenly assigned Joey’s cake to the \$125 cake. Additionally, the model might fail to explore alternative paths, leading to incorrect assessment of current states. Other mistakes include drawing a wrong conclusion despite correct deductions (Inconsistent Reasoning), failure to recognize conflicting information (Conflicts Resolving Failure), nonexistent constraints (False Premise), and a few instruction-following errors.
@@
4.5 State-Transition Performance Breakdown
@@
The right chart of Figure 4 shows the errors typically made by models during state transition. At solvable states, common errors include making multiple moves (Multiple Moves), violating Sudoku rules (Invalid Move), and transitioning to an unsolvable child state (Unsolvable Child). At unsolvable states, two primary errors are: failing to return to the parent state (Backtracking Failure) and making an additional move to a sibling state after backtracking (Sibling). Among these errors, Backtracking Failure is the most frequent across all models. Models sometimes jump back more than one level (e.g., to a grandparent state) or to a wrong state, indicating that LLMs struggle with step-by-step backtracking. For reasoning models (o1 and Gemini-FT), transitioning to siblings is the second most frequent error. This error is due to violating the minimal move principle (Table 1), highlighting weaknesses in their instruction-following capability. For general models, they frequently commit an invalid move. This shows that general models often fail to adequately check Sudoku rules.
@@ 4.6 Performance by Difficulty Level
To understand the state-checking performance across difficulty levels, we plot the density diagrams of correct vs. incorrect predictions by the number of unfilled cells in the current Sudoku state. We observe that Sudoku states with fewer empty cells are more likely to be predicted correctly. As the number of unfilled cells increases, the problem becomes more complex and requires more exploration, and the proportion of incorrect predictions increases.
@@
5 Training Results
As highlighted in §4, most models still lack state-checking and state-transition abilities, which are essential for reasoning. We hypothesize that training on these tasks can enhance performance on other reasoning-intensive tasks like math. To validate our hypothesis, we construct a training set consisting of state-checking and state-transition data from Sudoku, Graph Coloring, and Game of 24. We then train models on a mixture of this puzzle data and standard math data to test whether reasoning skills transfer beyond the puzzles themselves, ultimately improving overall mathematical reasoning.
Experimental setup.
The puzzle states are easily verified, making it suitable for Reinforcement Learning. Specifically, we perform GRPO (Shao et al., 2024) on DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B (DeepSeek-AI et al., 2025). We prepare $10,000$ samples from our puzzle dataset, and another $15,000$ samples from MetaMathQA (Yu et al., 2024), a popular training set for mathematical reasoning. Other training details are reported in Appendix A.3. Note that due to limited resources, we restrict the maximum completion length to 1024 in both training and evaluation.
Improvements on math benchmarks.
We start with $2,000$ training samples, with a $1:1$ distribution between math and puzzle data. We compare with two baselines: first is the base model, second is training with entirely math samples. The results in Table 5 show that combining puzzle data with MetaMath yields the highest accuracy on both GSM8K and MATH for both models, outperforming training on MetaMath alone. This consistent performance improvement suggests that the state-checking and state-transition logic of puzzle-solving generalize to more general mathematical problems, aligning well with our initial hypothesis.
@@ 5 Training Results
Analysis on the optimal ratio.
@@
To investigate the optimal ratio between puzzle-based and math-based data, we vary the proportion of math samples from $0.4$ to $1.0$ in a combined training set of 10k samples. In Figure 6, the performance on GSM8K steadily improves as the math ratio increases, peaking at a ratio of $0.8$. Beyond this point, increasing the math ratio further results in lower accuracy. This suggests that neither pure math training nor pure puzzle training is optimal. Instead, a balanced combination of puzzle-based and traditional math data provides the best performance. This indicates that our puzzle-based data though simple can complement the reasoning in standard math problems.
Analysis on the scaling effect.
@@
We investigate the scaling effect of using math-only and mixed math-puzzle data. We keep the ratio of math to $0.8$ in the mixed data. The results in Figure 7 show that, as we increase the number of training samples, both approaches benefit from scaling up. Noticeably, the mixed approach consistently outperforms math-only training starting from $5k$ samples. While math-only training shows diminishing returns or even a slight decline beyond $7.5k$ samples, the mixed approach continues to improve, reaching an accuracy peak of $81.3\%$ with around $12.5k$ samples. This scaling effect suggests the great potential of our simple puzzle data for enhancing the overall reasoning capability of LLMs.
00.40.50.60.70.80.91707070707272727274747474767676767878787880808080Ratio = Math / TotalAccuracy (%)
Figure 6: The effect of training data mixing (Math v.s. Puzzle) on GSM8K using the 1.5B model.
2.557.51012.5157474747476767676787878788080808082828282Number of training samples (k)Accuracy (%)Math-onlyMixed
Figure 7: The scaling performance on GSM8K using the 1.5B model.
@@ 6 Related Work
LLM Reasoning.
Advancing the reasoning capabilities of large language models is a critical goal in natural language processing Wos et al. (1992); Yang et al. (2018). In recent years, LLMs, combined with prompting techniques such as Chain of Thought Wei et al. (2022), Tree of Thought Yao et al. (2023), and Self-Consistency Wang et al. (2023), have shown remarkable performance in various reasoning tasks Cobbe et al. (2021); Srivastava et al. (2022). Current evaluation methods focus mainly on the final accuracy in reasoning-intensive domains, including mathematics Cobbe et al. (2021); Hendrycks et al. (2021); Chen et al. (2023); Rein et al. (2023); Ma et al. (2024), coding Chen et al. (2021); Austin et al. (2021), commonsense Mihaylov et al. (2018); Hendrycks et al. (2020), and logical reasoning Yao et al. (2023); Long (2023). However, as inference-time scaling gains importance Snell et al. (2024); Guo et al. (2025) and models are becoming more capable of reasoning, it is crucial to assess how effectively models perform reflection and correction during reasoning. While Tyagi et al. (2024) manually analyze the reasoning chains in logic puzzles, their approach lacks scalability. Some studies Singh et al. (2024); Zeng et al. (2024) evaluate how models handle reasoning mistakes, but these investigations often rely on rule-based mistakes that may be easily resolved by current LLMs. Moreover, these studies only assess reflection on past steps in a static manner. In our work, we address these limitations by introducing two novel tasks designed to more accurately reflect models’ capabilities in dynamic reasoning and error correction.
@@
Logic puzzles, which require deducing solutions from a set of rules Giadikiaroglou et al. (2024), are ideal for evaluating LLMs’ reasoning capabilities as they rely minimally on prior knowledge Li et al. (2024). Recent studies have explored LLMs on various puzzles with different emphases Mittal et al. (2024), such as Sudoku Ishay et al. (2023); Long (2023) for strategic thinking, Game of 24  Ding et al. (2023); Yao et al. (2023) for arithmetic calculations. Some investigate grid puzzles Dziri et al. (2024); Tyagi et al. (2024), crosswords Yao et al. (2023), chess puzzles Feng et al. (2024), mazes Noever and Burdick (2021), and Minesweeper Li et al. (2024). However, evaluation remains primarily focused on final accuracy.
@@ 7 Conclusion
In this work, we introduced FineReason, a novel logic-puzzle benchmark designed to comprehensively evaluate the reasoning capabilities of LLMs. Unlike existing benchmarks that focus mainly on final-answer accuracy, FineReason delves into intermediate reasoning steps, specifically emphasizing state checking and transition actions. This fine-grained evaluation captures a model’s ability to reflect, lookahead, and backtrack, which are vital aspects of human-like System 2 reasoning. Our experiments reveal significant gaps between reasoning-oriented and general-purpose LLMs, emphasizing the need to consider reflection and correction for robust reasoning evaluation. Furthermore, using puzzle-based data for training can enhance performance in broader mathematical tasks, highlighting the scalability of this approach and its potential to complement reasoning in standard math problems.
@@
In our study, we use textual tables to represent puzzle states. Our evaluation shows that models can reasonably understand this table format. However, there is potential to explore alternative representation formats, such as coordinates or images. The image format could be particularly valuable for facilitating the evaluation of multi-modal reasoning, offering a promising direction for future extensions of our work.
We adopt zero-shot CoT Kojima et al. (2022) as the prompt format in all our evaluations. While advanced prompting techniques have the potential to enhance performance, their use might shift the focus away from evaluating the genuine reasoning capabilities of LLMs without excessive reliance on external assistance. By focusing on zero-shot CoT, we aim to provide a clearer evaluation of the models’ inherent reasoning abilities.
@@ A.1 Dataset Statistics
Table 6 presents the statistics for the four tasks, including the total number of questions, as well as the number of solvable and unsolvable states for each task. For grid puzzles, we can only sample 94 solvable states with unsolvable children, resulting in a somewhat imbalanced dataset. Nonetheless, we have maintained balance between solvable and unsolvable states for the remaining three puzzles.
@@ A.2 Prompt Templates
Table 9 shows the state checking and state transition prompts for a Sudoku example.
A.3 Training Details
We train our models using GRPO based on OpenR1222https://github.com/huggingface/open-r1. We use one GPU to run vLLM for faster generation and the remaining GPUs for training. The hyperparameters and training details are reported in Table 7.
Learning rate 4e-5
Warm up ratio 0.1
Batch size 112
Max prompt length 1024
Max completion length 1024
Training epochs 1
Hardware 8 H20 (120 GB)
Table 7: Hyperparameter and training details.
A.4 Additional Experimental Results
@@
State Checking
You are given a partially filled 9x9 Sudoku grid represented as a list of lists, where empty cells are represented as 0. Your task is to determine if this current state can lead to a solvable solution. Specifically, use lookahead techniques to determine if it’s possible to fill the remaining cells according to standard Sudoku rules, ensuring that each row, column, and 3x3 subgrid contains unique numbers from 1 to 9.

### v2

@@ 4.4 Quality Analysis of State Checking
To examine the errors made in state checking, we conduct a human analysis of the mistakes from the best-performing model, o1, in Figure 3. The most common error is the Misinterpretation of Premises, where o1 incorrectly uses available information to reach a faulty conclusion. For instance, in a grid puzzle, given the clue “The chocolate piece, Joey’s cake, and the \$125 cake are three different cakes,” the model still mistakenly assigned Joey’s cake to the \$125 cake. Additionally, the model might fail to explore alternative paths, leading to an incorrect assessment of current states. Other mistakes include drawing a wrong conclusion despite correct deductions (Inconsistent Reasoning), failure to recognize conflicting information (Conflicts Resolving Failure), nonexistent constraints (False Premise), and a few instruction-following errors. Examples of the errors are shown in Appendix A.4.
@@
4.5 Model Behaviors in State Transition
@@
The right chart of Figure 4 shows the errors typically made by models during state transition. At solvable states, common errors include making multiple moves (Multiple Moves), violating Sudoku rules (Invalid Move), and transitioning to an unsolvable child state (Unsolvable Child). At unsolvable states, two primary errors are: failing to return to the parent state (Backtracking Failure) and making an additional move to a sibling state after backtracking (Sibling). Examples of the errors are shown in Appendix A.5. Among these errors, Backtracking Failure is the most frequent across all models. Models sometimes jump back more than one level (e.g., to a grandparent state) or to a wrong state, indicating that LLMs struggle with step-by-step backtracking. For reasoning models (o1 and Gemini-FT), transitioning to siblings is the second most frequent error. This error is due to violating the minimal move principle (Table 1), highlighting weaknesses in their instruction-following capability. For general models, they frequently commit an invalid move. This shows that general models often fail to adequately check Sudoku rules.
@@ 4.6 Performance by Difficulty Level
To understand the state-checking performance across difficulty levels, we plot the density diagrams of correct vs. incorrect predictions by the number of unfilled cells in Sudoku states (Figure 5). We find that Sudoku states with fewer empty cells are more likely to be predicted correctly. As the number of unfilled cells increases, the problem becomes more complex and requires more exploration, and the proportion of incorrect predictions increases. This likely reflects the increased computational complexity of looking ahead and determining solvability when many possibilities exist.
@@
5 Training on puzzle data
As highlighted in §4, most models struggle with identifying dead ends and backtracking. These observations reveal critical bottlenecks in models’ reflection and correction abilities, which are essential for reasoning. We hypothesize that training on state-checking and state-transition tasks can provide synthetic reflection, allow models to practice error correction, and ultimately transfer these reasoning skills to mathematical problem solving.
To validate our hypothesis, we construct a training set consisting of state-checking and state-transition data from Sudoku, Graph Coloring, and Game of 24. We train models on a mixture of puzzle data and math data to test whether reasoning skills transfer beyond puzzles themselves, ultimately improving performance on math tasks.
5.1 Experimental setup
We prepare $10,000$ puzzle samples and another $15,000$ samples from MetaMathQA (Yu et al., 2024), a popular training set for mathematical reasoning. The puzzle states are easily verified, making it suitable for Reinforcement Learning. Specifically, we train GRPO (Shao et al., 2024) on DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B (DeepSeek-AI et al., 2025). Other training details are reported in Appendix A.6. Due to limited computing resources, we restrict the maximum completion length to 1024 in both training and evaluation.
5.2 Improvements on mathematical reasoning
We start with $2,000$ training samples, with a $1:1$ distribution between puzzle and math data. We compare with two baselines: the first is the base model, and the second is training with entirely math samples. The results in Table 5 show that combining puzzle data with math data yields the highest accuracy on both GSM8K and MATH for both models, outperforming training on math data alone. This consistent performance improvement suggests that the state-checking and state-transition logic of puzzle solving generalizes to mathematical problems, aligning well with our initial hypothesis.
@@ 5.2 Improvements on mathematical reasoning
To further assess the impact of the mixed training, we analyze the number of correctly solved samples at each difficulty level in the MATH dataset, where a higher level means higher difficulty. Results in Appendix A.7 indicate that mixed training is especially effective for harder problems (levels 4 and 5), likely because complex problems require more reflection and correction. By incorporating puzzle-solving data into math reasoning training, we encourage models to reflect and backtrack when necessary.
To investigate the nature of improvements, we examine corrected examples after training. Around 60% of these examples use longer reasoning steps, determined by sentence splitter. Through quality analysis of 20 examples, we find more evidence of verification and self-correction in the thought process. Examples are shown in Appendix A.8.
00.40.50.60.70.80.91707070707272727274747474767676767878787880808080Proportion of math samples ($r_{m}$)GSM8K Accuracy (%)
Figure 6: The effect of math ratio on the 1.5B model.
Data GSM8K
None 65.5
Math-only 73.6
Mixed ($r_{u}=0.2$) 74.8
Mixed ($r_{u}=0.5$) 76.1
Mixed ($r_{u}=0.8$) 77.4
Table 6: The effect of unsolvable ratio on the 1.5B model.
5.3 Effect of the puzzle ratio
@@
To study the optimal ratio between puzzle-based and math-based data, we vary the proportion of math samples, $r_{m}$, from $0.4$ to $1.0$ in a combined training set of $10k$ samples. In Figure 6, the performance on GSM8K steadily improves as the math ratio increases, peaking at a ratio of $0.8$. Beyond this point, increasing the math ratio further results in lower accuracy. This suggests that neither pure math training nor pure puzzle training is optimal. Instead, a balanced combination of puzzle-based and math data provides the best performance. This indicates that our puzzle-based data, though simple, can complement the reasoning in standard math problems.
In addition, our previous analysis shows that models struggle significantly more with unsolvable states than solvable states in both state checking (§4.3) and state transition (§4.5). Our error analysis also shows that “backtracking failure” is a major error (§4.4), even for top reasoning models like o1. This means we could prioritize the learning for backtracking from these unsolvable cases to maximize learning benefits. To validate this, we examine the effect of varying the ratio of unsolvable data, $r_{u}$, in mixed training. In Table 6, it is evident that increasing the proportion of unsolvable data boosts the performance on GSM8K, suggesting that focusing on these unsolvable cases enhances the model’s ability to detect dead ends and revise strategies—skills transferable to math problem-solving.
5.4 Analysis on the scaling effect
@@
We examine the scaling effect of using math-only and mixed data for training. We use the optimal math ratio ($r_{m}=0.8$) in the mixed data. The results in Figure 7 show that, as we increase the number of training samples, both approaches benefit from scaling up. Noticeably, the mixed approach consistently outperforms math-only training starting from $5k$ samples. While math-only training shows diminishing returns or even a slight decline beyond $7.5k$ samples, the mixed approach continues to improve, reaching an accuracy peak of $81.3\%$ with around $12.5k$ samples. This scaling effect suggests the great potential of our simple puzzle data for enhancing the overall reasoning capability of LLMs.
2.557.51012.5157474747476767676787878788080808082828282Number of training samples (k)GSM8K Accuracy (%)Math-onlyMixed
Figure 7: The scaling performance of the 1.5B model.
5.5 Improvements on general reasoning
To investigate if our puzzle data benefits general reasoning tasks beyond math, we conduct additional experiments on MMLU (Hendrycks et al., 2021a), a dataset covering 57 tasks including STEM, humanities, social sciences, and others. Using the same optimal ratio ($r_{m}=0.8$) as in math training, we observe consistent performance improvements over MMLU-only training (Table 7).
Next, we analyze the performance improvement across different subsets in MMLU. Results in Appendix A.9 show that mixed training significantly improves performance on complex, multi-step reasoning tasks that might require more reflection (STEM +10%, Social Sciences +3.8%), but shows less benefit for simpler, more direct question types.
Overall, our puzzle data bridges the gap of the lack of annotated intermediate training data for general reasoning tasks by providing synthetic, structured scenarios for learning reasoning skills. While real-world reasoning is less deterministic, our results indicate that mastering deterministic steps in puzzle solving builds foundational skills like reflection and correction, which demonstrably transfer to general reasoning tasks.
Model Data MMLU-test
DeepSeek-R1-Distill-Qwen-1.5B None 37.8
MMLU-train-only 41.1
Mixed 44.9
DeepSeek-R1-Distill-Qwen-7B None 54.9
MMLU-train-only 61.2
Mixed 64.2
Table 7: Training with our puzzle data improves general reasoning tasks on MMLU.
@@ 6 Related Work
Reasoning Capabilities of LLMs.
Advancing the reasoning capabilities of large language models is a critical goal in natural language processing Wos et al. (1992); Yang et al. (2018). In recent years, LLMs, combined with prompting techniques such as Chain of Thought Wei et al. (2022), Tree of Thought Yao et al. (2023), and Self-Consistency Wang et al. (2023), have shown remarkable performance in various reasoning tasks Cobbe et al. (2021); Srivastava et al. (2023). Current evaluation methods focus mainly on the final accuracy in reasoning-intensive domains, including mathematics Cobbe et al. (2021); Hendrycks et al. (2021b); Chen et al. (2023); Rein et al. (2024); Ma et al. (2024), science Mihaylov et al. (2018); Xu et al. (2021a, b); Huang et al. (2025, 2024), coding Chen et al. (2021); Austin et al. (2021), commonsense Hendrycks et al. (2021a), and logical reasoning Yao et al. (2023); Long (2023). However, as inference-time scaling gains importance Snell et al. (2025); DeepSeek-AI et al. (2025) and models become more capable of reasoning, it is crucial to assess how effectively models perform reflection and correction during reasoning. While Tyagi et al. (2024) manually analyzes the reasoning chains in logic puzzles, their approach lacks scalability. Some studies Singh et al. (2024); Zeng et al. (2024); Xu et al. (2024) evaluate how models handle reasoning mistakes, but these investigations often rely on rule-based mistakes that may be easily resolved by current LLMs. Moreover, these studies only assess reflection on past steps in a static manner. In our work, we address these limitations by introducing two novel tasks designed to more accurately reflect models’ capabilities in dynamic reasoning and error correction.
@@
Logic puzzles, which require deducing solutions from a set of rules Giadikiaroglou et al. (2024), are ideal for evaluating the reasoning abilities of LLMs, as they rely minimally on prior knowledge Li et al. (2024b). Recent studies have explored LLMs on various puzzles with different emphases Mittal et al. (2024), such as Sudoku Ishay et al. (2023); Long (2023) for strategic thinking, Game of 24 Ding et al. (2024); Yao et al. (2023) for arithmetic calculations. Some investigate grid puzzles Dziri et al. (2023); Tyagi et al. (2024), crosswords Yao et al. (2023), chess puzzles Feng et al. (2023), mazes Noever and Burdick (2021), Minesweeper Li et al. (2024b), and abstract pattern recognition Chia et al. (2024b). However, the evaluation remains mainly focused on the final accuracy.
@@ 7 Conclusion
In this work, we introduce FineReason, a novel logic-puzzle benchmark designed to comprehensively evaluate the reasoning capabilities of LLMs. Unlike existing benchmarks that focus mainly on final-answer accuracy, FineReason delves into intermediate reasoning steps, specifically emphasizing state checking and transition actions. This fine-grained evaluation captures the model’s ability to reflect, look ahead, and backtrack, which are vital aspects of human-like System 2 reasoning. Our experiments reveal significant gaps between reasoning-oriented and general-purpose LLMs, emphasizing the need to consider reflection and correction for robust reasoning evaluation. Furthermore, using puzzle-based data for training can enhance performance in general reasoning tasks, highlighting the scalability of this approach and its potential to transfer reasoning skills beyond puzzles to broader reasoning.
Acknowledgment
This research is supported by DAMO Academy through DAMO Academy Research Intern Program and Alibaba-NTU Singapore Joint Research Institute (JRI), Nanyang Technological University, Singapore. We would also like to extend our gratitude to Interdisciplinary Graduate Programme and College of Computing and Data Science of NTU, for their support.
@@ Acknowledgment
Our study has several limitations. First, we represent puzzle states using textual tables. Our evaluation shows that models can reasonably understand this table format. However, there is potential to explore alternative representation formats, such as coordinates or images. The image format could be particularly valuable for facilitating the evaluation of multi-modal reasoning, offering a promising direction for future extensions of our work. Second, we employ zero-shot CoT prompting Kojima et al. (2022) to focus on evaluating the inherent reasoning capabilities of LLMs, foregoing more advanced prompting techniques that could potentially improve performance but might obscure the models’ true reasoning abilities. Finally, our current evaluation prioritizes reasoning correctness over efficiency, treating all solvable states equally regardless of the steps required to reach the solution. Future research could incorporate metrics like "steps to solution" to assess efficiency.
@@ A.1 Dataset Statistics
Table 8 presents the statistics for the four tasks, including the total number of questions, as well as the number of solvable and unsolvable states for each task. For grid puzzles, we can only sample 94 solvable states with unsolvable children, resulting in a somewhat imbalanced dataset. Nonetheless, we have maintained a balance between solvable and unsolvable states for the remaining three puzzles.
@@ A.2 Prompt Templates
In Table 13,  14,  15,  16,  17, we show the state checking and state transition prompts for each puzzle.
A.3 Additional results on state checking precision, recall and F1 scores
@@
A.4 Error types in state checking
Through human analysis, we categorise five common types of errors o1 made in state checking. We show one example for each type of error: Misinterpretation of Premises (Table 18), Exploration and Backtracking Issues (Table 19), Inconsistent Reasoning (Table 20), Conflicts Resolving Failure (Table 21), and False Premise (Table 22).
A.5 Error types in state transition
We automatically classify the state transition mistakes based on the next move. We show one example for each type of error models made in the state transition: Multiple Moves (Table 23), Invalid Move (Table 24), Unsolvable Child (Table 25), Backtracking Failure (Table 26), Sibling (Table 27).
A.6 Training Details
We train our models using GRPO based on OpenR1333https://github.com/huggingface/open-r1. We use one GPU to run vLLM for faster generation and the remaining GPUs for training. The hyperparameters and training details are reported in Table 10.
Learning rate 4e-5
Warm up ratio 0.1
Batch size 112
Max prompt length 1024
Max completion length 1024
Training epochs 1
Hardware 8 GPUs (120 GB)
Table 10: Hyperparameter and training details.
A.7 Benefits of mixed-training across difficulty level
Table 11 analyses the impact of mixed training on performance across varying difficulty levels within the MATH dataset. Notably, the gains achieved through mixed training are most significant for levels 4 and 5 problems, which represent the most challenging instances. This suggests that the benefits of mixed training are amplified when tackling complex problems that necessitate more extensive reflection and iterative correction processes.
Difficulty Level Math-only Mixed Difference (%)
Level 1 372 386 4%
Level 2 622 630 1%
Level 3 675 661 -2%
Level 4 540 588 9%
Level 5 348 389 12%
Table 11: The % improvement of correct samples from mixed training using the 1.5B model.
A.8 Quality analysis of model outputs with/without mixed training
Through quality analysis of 20 examples, we found more evidence of verification and self-correction in their thought process. Here we show two examples in Table 28.
In the first example, model 1 applied the formula directly without cross-checking its consistency with the given example of the third row. After training, model 2 verified the calculation against the provided example before reaching the answer.
In the second example, model 2 initially included an incorrect letter “C” but it caught this mistake and removed “C” before proceeding. However, model 1 skipped this step.
These examples suggest that puzzle training helps reinforce skills like constraint validation and error correction, which generalize to math reasoning.
A.9 Benefits of mixed-training across different reasoning tasks
In Table 12, we analyze the performance improvement across different subsets in MMLU. We organize the tasks into four categories following  Hendrycks et al. (2021a). The most significant improvement is in STEM tasks, where accuracy increases by 10%. The STEM subset consists of multi-step reasoning problems that require reflection and correction in the reasoning process. We also observe some benefits (+3.8%) to the social sciences subset, as some subjects (eg, Econometrics) require multi-step reasoning. The humanities subset and other tasks, which are typically straightforward questions requiring fewer reasoning steps, show smaller improvements.
Category MMLU-only Mixed Improvement (%)
Humanities 33.0 33.8 0.8
Social Sciences 45.4 49.2 3.8
STEM 48.7 58.7 10.0
Others 41.2 43.3 2.1
Table 12: The improvement in MMLU performance from mixed training using the 1.5B model across different task categories.
Sudoku State Checking
You are given a partially filled 9x9 Sudoku grid represented as a list of lists, where empty cells are represented as 0.
Your task is to determine if this current state can lead to a solvable solution. Specifically, use lookahead techniques to determine if it’s possible to fill the remaining cells according to standard Sudoku rules, ensuring that each row, column, and 3x3 subgrid contains unique numbers from 1 to 9.